A statistics Department had purchased 24calculators in which 4 are defective. Calculators are selected one-after-another without replacement and tested. What is the probability that the second calculator found to be defective is the eighth calculator selected.

1
Expert's answer
2022-01-04T08:33:14-0500

Use the Hypergeometric distribution

The probability with "n=7, N=24, k=4"


"P(X=1)=\dfrac{\dbinom{4}{1}\dbinom{24-4}{7-1}}{\dbinom{24}{7}}=\dfrac{340}{759}"

"P(second\ defective)=\dfrac{\dbinom{4}{1}\dbinom{24-4}{7-1}}{\dbinom{24}{7}}\cdot\dfrac{4-1}{24-7}"

"=\dfrac{340}{759}\cdot\dfrac{3}{17}=\dfrac{20}{253}"

The probability that the second calculator found to be defective is the eighth calculator selected is "\dfrac{20}{253}."